# AuscultaBase: A Foundational Step Towards AI-Powered Body Sound Diagnostics

### v1

@@ 1 Introduction
  • AuscultaBase-Bench: A Comprehensive Benchmark for Auscultation Evaluation. To facilitate rigorous assessment, AuscultaBase-Bench provides a structured evaluation framework across three primary diagnostic tasks—abnormality detection, disease classification, and activity recognition—spanning 16 sub-tasks that reflect the diverse and complex demands of clinical auscultation. This benchmark enables consistent, standardized performance comparisons for current and future auscultation models.
2 Related Works
The auscultatory body sounds include heart sounds, respiratory sounds, and bowel sounds. Existing models typically focus on one specific type of body sound. For heart sounds, the primary tasks concentrate on abnormal detection and disease diagnosis. Most methods employ traditional acoustic features for feature processing. For example, Chen et al. (2023) initially applied noise reduction to the audio and then utilized Short-Time Fourier Transform (STFT) to obtain the audio spectrogram. Their main network was constructed using a CNN and included an attention module at the end. Guo et al. (2023) combined high-order spectral estimation with STFT for audio feature extraction and designed a dual-stream CNN as the primary network. Some approaches have also introduced general audio pre-training models, such as the study that compared the effectiveness of features extracted using PANN (Kong et al., 2020) with those derived from non-pre-trained models using STFT features alongside architectures like VGG and ResNet, demonstrating the transferable capabilities of general audio pre-training models in the field of heart sound diagnosis (Koike et al., 2020).
For respiratory sounds, the tasks are similarly focused on abnormal detection and disease diagnosis. Early methods employed traditional feature extraction techniques such as STFT and MFCC (Oletic & Bilas, 2017; Jung et al., 2021). In recent years, due to the COVID-19 pandemic, several pre-training models targeting respiratory sounds have emerged. Zhang et al. (2024) utilized cough sound data collected via microphones and respiratory sound data obtained from stethoscopes to construct a pre-training dataset, comparing two contrastive methods and one generative method, demonstrating the efficacy of the pre-training model across 19 tasks involving cough and respiratory sounds. Baur et al. (2024) developed a health acoustic event detector that identifies audio clips from YouTube corresponding to specific events, including coughing, breathing, throat clearing, laughing, and speaking, followed by pre-training using a generative approach based on MAE (He et al., 2022), and subsequently fine-tuning on downstream tasks related to the identified events.
Regarding bowel sounds, there is a relative lack of datasets and methods, with the primary task being bowel sound detection. Among them, most methods have been tested on proprietary datasets. Sitaula et al. (2022) first collected bowel sound data from 49 newborns and constructed a model combining CNN with Laplace’s hidden semi-markov model after extracting audio features using MFCC. Baronetto et al. (2023) collected bowel sound data from 18 participants and explored the effects of using Audioset for pre-training, along with various CNN and attention-based architectures.
Existing methods overlooked the interrelatedness of various body sounds in auscultation, focusing exclusively on individual categories of sounds. In contrast, the proposed A4Body-Model fully considers this correlation by utilizing all data related to auscultation for pre-training. In this way, our approach meets the high data requirements for pre-training while eliminating the potential noise impact from unrelated audio.
3 System Overview
The system overview of AuscultaBase is exhibited in Figure 1, which consists of three main components: pretraining datasets (-Corpus), foundation model(-Model), and evaluation benchmark (-Bench). Firstly, the -Corpus is curated by the collection of existing body sound datasets and then processed (Section 4) for subsequent pretraining the -Model as elaborated in Section 5. Then the -Bench is introduced to evaluate existing models comprehensively including 16 tasks in Section 6.
4 AuscultaBase-Corpus: Pretraining Datasets
Figure 1: System overview of AuscultaBase. The AuscultaBase-Corpus is first collected by (a) Audio Collection and then processed by (b) Data Processing. AuscultaBase-Model is obtained by (c) Foundation Model Pretraining and evaluated with AuscultaBase-Bench: (d) Benchmark. Based on the tasks, the fine-tuned model can be applied for various scenarios of (e) Clinical Auscultation.
11 datasets are collected to curate AuscultaBase-Corpus and enable the pretraining of our auscultation model, which involves lung, heart, and bowel sounds (shown in Table 1). After filtering the broken audio clips, the entire pertaining datasets consist of 40, 317 samples, with a total duration of 322.4 hours. All of them can be publicly available except for HSDreport and XHheartSound, in which XHheartSound is a private dataset collected by Xinhua Hospital Affiliated To Shanghai Jiao Tong University School of Medicine and has not been open source yet. All of them were made with different versions of digital stethoscopes attached to the chest, heart, and bowel respectively. This allows our model to acquire more powerful capabilities for identifying the frequency and beat of the body sounds, and then better boosting the application to clinical auscultation. It is noted that some of the labeled samples are held out from the pertaining datasets, which can be used for subsequent downstream benchmarking.
Table 1 demonstrates the statistics of our pretraining datasets. As the data sources vary such as in the auscultation position, sampling rate, and durations, we first resample each recording to 16kHz and merge them into a mono channel. After that, the beginning and ending silence is removed to ensure the quality of the recordings. The trimmed clips are then transformed using 64 LogMel Spectrogram with a Hanning window which shits every 32 ms. The spectrograms are then standardized with Min-Max Normalization and finally used to pretrain our auscultation foundation model.

### v2

@@ 1 Introduction
  • AuscultaBench: A comprehensive evaluation benchmark to assess AI-based auscultation models. To facilitate rigorous assessment, AuscultaBench provides a structured evaluation framework across two primary diagnostic tasks—abnormality detection and disease classification—spanning 16 sub-tasks that reflect the diverse and complex demands of clinical auscultation. This benchmark ensures consistent and reproducible performance comparisons, facilitating advancements in automated diagnostic methodologies.
Figure 1: System overview of our framework. During the Data Construction stage, the AuscultaCorpus is first collected by stethoscopes and then pre-processed into spectrograms. Followed by the AuscultaCorpus, the audio encoder is pre-trained with contrastive learning to obtain the base model for auscultation, AuscultaBase. Finally, we apply linear/full fine-tuning to AuscultaBase to adapt to 16 downstream auscultation tasks within our evaluation benchmark, AuscultaBench. Additionally, we conduct clinical evaluations between our model and experienced clinicians to analyze the auscultation performance, extending the prospect for various scenarios of clinical applications.
2 Results
2.1 Overview of proposed datasets & model & benchmark
The system overview of our framework is illustrated in Figure 1, comprising three key components: the pretraining dataset (AuscultaCorpus), the foundation model (AuscultaBase), and the evaluation benchmark (AuscultaBench). AuscultaCorpus is constructed by aggregating diverse body sound datasets and preprocessing them for subsequent model pretraining.
AuscultaCorpus is curated from 11 datasets to enable the pretraining of our auscultation model, covering lung, heart, and bowel sounds. After filtering out corrupted audio clips, the final pretraining dataset comprises 40,317 samples with a total duration of 322.4 hours. The statistics of these datasets are summarized in Extended Table 2.
We pretrain our auscultation model using the constructed datasets to build our auscultation foundation model: AuscultaBase. For datasets that already include distinct training and validation splits, we preserve the provided validation sets. For those without an explicit division, we randomly select 10% of the samples for validation after shuffling the data.
To evaluate the auscultation capabilities of current foundation models, we introduce AuscultaBench—a comprehensive benchmark derived from the datasets collated. The statistics of AuscultaBench is exhibited in Figure 2, and a more specific distribution is detailed in Extended Table 2. Importantly, all test sets originally associated with these datasets were excluded beforehand to prevent any data leakage.
2.2 Overall Performance Across Diverse Tasks
Figure 2: Statistics of AuscultaBench. The Sankey diagram shows how the different body sound types (left), datasets (left middle), tasks (right middle), and abnormality or disease categories (right) contribute to the final evaluation benchmark. On the left of the bottom, two bar charts show the data distributions on cardiac and respiratory entities respectively.
Tables 1 summarize the Macro-F1 and Micro-F1, Precision, and Recall scores, respectively. Our proposed model, AuscultaBase, demonstrates clear and consistent improvements over the baseline models across these varied tasks. Notably, AuscultaBase achieved the highest Macro-F1 performance in 12 out of 16 tasks and led in Micro-F1 for 9 tasks. By comparison, the baselines—such as OPERA-CT, AudioMAE, CLAP, and PANN—led in at most a single task each. For instance, PANN led only in tasks T6 and T10 for Micro-F1, and even then, its performance did not translate to improvements in Macro-F1. This consistent superiority across different body sound categories (lung, heart, and bowel) underscores the enhanced capability of AuscultaBase to extract distinctive and generalizable acoustic features, which are critical for accurate and reliable body sound classification.
Balanced Accuracy Across Imbalanced Datasets
A further comparison between Macro-F1 and Micro-F1 scores reveals that AuscultaBase maintains balanced accuracy across classes, an essential quality given the inherent class imbalances in many medical datasets. In 9 out of 16 tasks (T1, T3, T7, T9, T11–T16), AuscultaBase simultaneously achieved the highest scores for both metrics. This balanced performance is particularly significant in clinical contexts, where ensuring that minority classes (often representing critical or rare conditions) are accurately identified is as important as the overall performance.
Task ID Function Sound Type Task Type OPERA-CT AudioMAE CLAP PANN AuscultaBase
Macro-F1
T1 AD L MC 44.99$\pm$1.53 38.14$\pm$1.24 47.71$\pm$0.09 25.99$\pm$0.07 48.14$\pm$0.46
T2 AD L MC 25.09$\pm$1.04 18.20$\pm$0.81 36.55$\pm$0.41 15.55$\pm$0.00 30.25$\pm$1.10
T3 AD L BC 54.49$\pm$1.36 55.72$\pm$0.11 55.63$\pm$0.63 40.99$\pm$0.00 57.28$\pm$0.78
T4 AD L MC 32.08$\pm$0.05 27.70$\pm$0.14 33.19$\pm$0.03 16.75$\pm$0.03 33.48$\pm$0.12
T5 AD L MC 22.04$\pm$0.08 14.42$\pm$0.00 23.24$\pm$0.65 11.29$\pm$0.15 23.74$\pm$0.01
T6 AD L MC 28.65$\pm$0.87 19.99$\pm$1.33 37.90$\pm$0.26 34.71$\pm$0.31 26.23$\pm$0.35
T7 AD L MC 23.95$\pm$0.13 9.34$\pm$1.53 27.53$\pm$2.91 7.80$\pm$0.00 28.03$\pm$3.66
T8 AD H MC 44.40$\pm$0.52 42.93$\pm$2.16 48.33$\pm$3.18 28.71$\pm$0.00 41.52$\pm$2.45
T9 AD B BC 69.82$\pm$1.38 45.60$\pm$0.85 71.99$\pm$1.02 66.40$\pm$1.29 72.88$\pm$0.05
T10 DD L MC 23.28$\pm$4.08 22.10$\pm$2.97 34.37$\pm$3.83 21.46$\pm$2.03 24.68$\pm$0.83
T11 DD L ML 15.10$\pm$7.12 7.31$\pm$4.18 7.99$\pm$0.07 0.00$\pm$0.00 20.10$\pm$5.85
T12 DD L MC 30.74$\pm$0.77 17.00$\pm$5.17 27.68$\pm$2.07 30.15$\pm$5.57 41.78$\pm$1.83
T13 DD H MC 77.37$\pm$1.94 50.18$\pm$15.23 86.82$\pm$0.32 84.19$\pm$2.15 88.10$\pm$0.69
T14 DD H BC 65.04$\pm$0.56 72.35$\pm$0.62 61.97$\pm$2.78 66.01$\pm$0.34 77.84$\pm$0.23
T15 DD H BC 58.10$\pm$1.14 61.66$\pm$0.45 59.94$\pm$1.12 48.38$\pm$0.37 62.84$\pm$1.49
T16 DD H BC 47.27$\pm$0.00 47.27$\pm$0.00 47.27$\pm$0.00 47.27$\pm$0.00 52.53$\pm$1.15
Micro-F1
T1 AD L MC 64.24$\pm$1.81 61.68$\pm$4.55 72.09$\pm$1.68 63.64$\pm$0.03 71.28$\pm$0.37
T2 AD L MC 64.02$\pm$0.22 63.21$\pm$0.16 61.03$\pm$0.97 63.61$\pm$0.00 65.73$\pm$0.12
T3 AD L BC 71.24$\pm$0.34 63.58$\pm$0.44 66.22$\pm$1.34 69.45$\pm$0.00 70.67$\pm$0.17
T4 AD L MC 51.32$\pm$0.29 46.54$\pm$0.19 44.84$\pm$0.35 49.60$\pm$0.01 50.46$\pm$0.21
T5 AD L MC 51.03$\pm$0.31 48.45$\pm$0.00 44.50$\pm$0.54 49.59$\pm$0.02 49.42$\pm$1.07
T6 AD L MC 46.77$\pm$2.54 32.98$\pm$5.55 53.86$\pm$0.38 54.86$\pm$0.15 41.75$\pm$0.82
T7 AD L MC 48.48$\pm$0.00 25.76$\pm$1.52 46.97$\pm$1.52 24.24$\pm$0.00 51.52$\pm$3.03
T8 AD H MC 77.53$\pm$0.00 80.06$\pm$0.32 81.33$\pm$0.32 75.63$\pm$0.00 78.16$\pm$0.32
T9 AD B BC 82.24$\pm$0.93 81.15$\pm$0.16 82.87$\pm$0.62 82.55$\pm$0.31 85.05$\pm$0.31
T10 DD L MC 90.81$\pm$0.26 91.34$\pm$0.26 91.73$\pm$0.66 92.13$\pm$0.00 90.81$\pm$1.31
T11 DD L ML 18.49$\pm$8.78 22.52$\pm$8.88 19.22$\pm$0.30 0.00$\pm$0.00 43.15$\pm$1.29
T12 DD L MC 36.00$\pm$2.00 43.00$\pm$1.00 41.00$\pm$3.00 40.00$\pm$4.00 49.00$\pm$1.00
T13 DD H MC 77.00$\pm$2.00 56.50$\pm$9.50 87.50$\pm$0.50 84.00$\pm$2.00 87.50$\pm$0.50
T14 DD H BC 81.83$\pm$0.47 81.52$\pm$0.47 80.90$\pm$0.16 76.86$\pm$0.31 87.27$\pm$0.31
T15 DD H BC 59.02$\pm$2.06 62.03$\pm$0.63 60.92$\pm$1.11 55.70$\pm$0.32 63.45$\pm$1.42
T16 DD H BC 89.66$\pm$0.00 89.66$\pm$0.00 89.66$\pm$0.00 89.66$\pm$0.00 90.52$\pm$0.65
Precision
T1 AD L MC 50.80$\pm$2.60 48.00$\pm$7.06 48.80$\pm$1.01 29.55$\pm$8.35 50.86$\pm$0.31
T2 AD L MC 39.39$\pm$8.63 24.73$\pm$0.18 38.52$\pm$1.75 12.72$\pm$0.00 38.77$\pm$0.73
T3 AD L BC 65.93$\pm$1.51 56.00$\pm$0.06 57.57$\pm$0.68 34.73$\pm$0.00 63.92$\pm$0.41
T4 AD L MC 40.11$\pm$0.39 32.97$\pm$0.35 35.55$\pm$0.18 18.60$\pm$0.07 42.11$\pm$0.08
T5 AD L MC 33.04$\pm$0.93 26.00$\pm$0.00 28.77$\pm$0.36 12.68$\pm$0.24 33.33$\pm$0.14
T6 AD L MC 32.36$\pm$5.23 24.93$\pm$2.36 38.42$\pm$0.20 36.38$\pm$2.44 30.31$\pm$0.13
T7 AD L MC 19.84$\pm$0.16 10.00$\pm$5.16 38.13$\pm$0.87 4.85$\pm$0.00 28.42$\pm$6.25
T8 AD H MC 64.57$\pm$0.44 54.66$\pm$2.81 71.40$\pm$17.25 25.21$\pm$0.00 56.75$\pm$1.96
T9 AD B BC 70.88$\pm$1.63 65.56$\pm$25.06 72.14$\pm$1.02 71.27$\pm$0.71 76.50$\pm$0.92
T10 DD L MC 22.81$\pm$4.26 21.50$\pm$3.18 45.01$\pm$1.54 32.64$\pm$5.97 26.16$\pm$0.73
T11 DD L ML 12.56$\pm$7.64 5.41$\pm$2.00 10.91$\pm$0.00 0.00$\pm$0.00 30.00$\pm$7.50
T12 DD L MC 40.90$\pm$3.90 13.57$\pm$5.18 33.68$\pm$1.26 34.71$\pm$11.37 38.78$\pm$0.28
T13 DD H MC 77.59$\pm$1.51 58.90$\pm$15.83 87.38$\pm$0.43 84.08$\pm$2.34 89.65$\pm$1.04
T14 DD H BC 74.70$\pm$2.18 72.47$\pm$0.50 72.60$\pm$0.79 65.66$\pm$0.42 84.55$\pm$1.22
T15 DD H BC 59.42$\pm$2.17 61.90$\pm$0.66 60.91$\pm$1.19 57.21$\pm$0.61 63.41$\pm$1.46
T16 DD H BC 44.83$\pm$0.00 44.83$\pm$0.00 44.83$\pm$0.00 44.83$\pm$0.00 78.41$\pm$16.61
Recall
T1 AD L MC 44.12$\pm$1.24 39.71$\pm$1.33 48.31$\pm$0.23 33.36$\pm$0.03 47.40$\pm$0.52
T2 AD L MC 25.82$\pm$0.82 21.10$\pm$0.52 38.89$\pm$0.47 20.00$\pm$0.00 29.03$\pm$0.85
T3 AD L BC 56.13$\pm$1.35 55.62$\pm$0.13 55.74$\pm$0.35 50.00$\pm$0.00 57.63$\pm$0.48
T4 AD L MC 32.80$\pm$0.13 30.06$\pm$0.84 33.74$\pm$0.00 25.05$\pm$0.00 33.43$\pm$0.14
T5 AD L MC 22.60$\pm$0.10 17.71$\pm$0.00 23.85$\pm$0.73 16.75$\pm$0.05 23.29$\pm$0.17
T6 AD L MC 28.89$\pm$0.32 25.75$\pm$2.61 38.89$\pm$0.03 35.47$\pm$0.02 28.47$\pm$0.23
T7 AD L MC 30.87$\pm$0.48 20.77$\pm$0.77 30.09$\pm$3.31 20.00$\pm$0.00 34.05$\pm$1.74
T8 AD H MC 42.02$\pm$0.36 41.91$\pm$1.75 45.32$\pm$1.74 33.33$\pm$0.00 40.52$\pm$1.64
T9 AD B BC 68.96$\pm$1.20 50.41$\pm$0.41 71.86$\pm$1.01 64.23$\pm$1.35 70.69$\pm$0.44
T10 DD L MC 25.92$\pm$6.03 22.74$\pm$2.74 30.01$\pm$3.78 20.49$\pm$1.86 24.97$\pm$0.75
T11 DD L ML 28.22$\pm$0.72 20.19$\pm$17.31 8.59$\pm$2.27 0.00$\pm$0.00 16.55$\pm$5.29
T12 DD L MC 43.17$\pm$0.32 25.00$\pm$5.00 31.27$\pm$2.14 33.85$\pm$4.72 52.62$\pm$1.67
T13 DD H MC 79.72$\pm$1.66 56.68$\pm$14.33 86.86$\pm$0.36 84.70$\pm$1.69 87.48$\pm$0.59
T14 DD H BC 62.88$\pm$0.54 72.74$\pm$1.25 60.43$\pm$2.25 66.73$\pm$0.54 74.43$\pm$0.07
T15 DD H BC 58.64$\pm$1.42 61.68$\pm$0.49 60.22$\pm$1.10 53.73$\pm$0.32 62.92$\pm$1.46
T16 DD H BC 50.00$\pm$0.00 50.00$\pm$0.00 50.00$\pm$0.00 50.00$\pm$0.00 52.64$\pm$0.56
Table 1: Macro-F1, Micro-F1, Precision, and Recall scores for abnormality detection (AD: T1-T8) and disease diagnosis (DD: T10-T16) tasks. The best model for each task is in bold. We report mean and standard deviation with 3 runs. Macro-F1 ranges from 0 to 100, and higher is better.
2.3 Binary Classification: Discriminative Power and AUC Analysis
To further probe the model’s discriminative capabilities, we computed the Area Under the Receiver Operating Characteristic curve (AUC) for select binary classification tasks (T3, T9, T14, T15, and T16). Unlike F1 scores, AUC provides an aggregate measure of performance across all classification thresholds. As shown in Figure 3, AuscultaBase consistently outperformed all baseline models, particularly in task T14, where it achieved an AUC of 0.920 compared to 0.872 for AudioMAE (the best baseline). This superior discriminative performance is critical for clinical applications, as it indicates a robust ability to distinguish between healthy and abnormal states, thus reducing the risk of misdiagnosis.
Figure 3: Receiver operating characteristic (ROC) curve and average AUC of binary classification tasks (T4, T9, T14, T15, and T16) with 3 independent runs (labeled with 3 colors).
Figure 4: Borda count scores (BCS) categorized by the task function (abnormality detection and disease diagnosis), sound type (lung sound, heart sound, and bowel sound), and task type (binary classification, multiclass classification, and multilabel classification). BCS ranges from 1 to 5, and higher is better.
2.4 Multidimensional Evaluation of Model Robustness
To provide a nuanced evaluation of model performance across multiple dimensions, we categorized the tasks by function, sound type, and task type, and computed the Borda Count Scores (BCS) based on the reported Macro-F1, Micro-F1, Precision, Recall, and AUC performance as depicted in Figure 4.
Performance by Function: Abnormality Detection and Disease Diagnosis
When tasks were grouped into abnormality detection and disease diagnosis, AuscultaBase always achieved the highest BCS in F1 scores and AUC. In disease diagnosis tasks, for example, AuscultaBase reaches a BCS of 4.86 (Macro-F1), a substantial improvement over competing models. This result not only highlights the model’s capacity for nuanced clinical assessment but also its potential to impact patient care through more precise disease identification.
Adaptability Across Sound Types
Performance was also evaluated based on the sound type, using BCS for lung, heart, and bowel sounds. As depicted in Figure 4, AuscultaBase consistently outperformed baseline models, particularly excelling in heart sound classification. This robust performance across diverse acoustic environments demonstrates the model’s adaptability, making it a versatile tool for real-world clinical applications where multiple sound sources and conditions are encountered.
Robustness Across Task Types
Finally, we assessed performance across various task formulations, including binary classification, multi-class classification, and multi-label classification. As shown in the right panel of Figure 4, AuscultaBase achieved top performance across all task types, with particularly strong results in binary classification tasks. This versatility indicates that the model is capable of effectively generalizing to different problem formulations, an important quality for its application in diverse diagnostic scenarios.
2.5 Comparative Analysis: AuscultaBase vs. Human Experts
2.5.1 Clinical Evaluation Setup
We further assessed the diagnostic performance of AuscultaBase by comparing its predictions with those of experienced human clinicians in detecting congenital heart disease (CHD). For this evaluation, we collected 100 heart sound recordings. The auscultation protocol consists of recordings over 5 body sites: aortic region (right 2nd intercostal space), pulmonic region (left 2nd intercostal space, parasternal), Erb’s point (left 3rd intercostal space aka left lower sternal border), tricuspid region (left 4th intercostal space, parasternal), mitral region (left 5th intercostal space, midclavicular). The auscultation positions are visualized in Figure 5. To detect sufficient cardiac cycles, at least 15s of heart sound using direct skin contact is obtained per site. The electronically amplified stethoscope (Littmann 3200, 3M) is used for data acquisition. During the examination the participant is seated, laid down, or held to the most comfortable position. Additionally, each recording is paired with a cardiac ultrasound report serving as the diagnostic gold standard. The dataset comprises recordings from 50 healthy children and 50 children diagnosed with CHD (including atrial and ventricular septal defects), with participant ages ranging from 11 days to 12 years (Figure 5). Five clinicians from Anonymous Hospital, each with over 15 years of experience, independently diagnosed these recordings without access to the corresponding ultrasound reports. In parallel, the AuscultaBase model was fine-tuned on the HSDReport dataset and evaluated on these same 100 samples, ensuring that the test recordings were excluded from the fine-tuning process.
Figure 5: The statistics of test samples and diagnostic performance comparison between human and AuscultaBase. (a) The overview of the clinical evaluation process. (b) The distribution of test samples across gender and age. (c) The confusion matrices of the diagnostic results derived by the human and AuscultaBase. (d) The sensitivity, specificity, and diagnostic accuracy across different genders and age groups. The sensitivity, specificity, and accuracy range from 0 to 1, and higher is better.
2.5.2 Diagnostic Performance: Insights and Implications
The results, depicted in Figures 5 and 5, yield several key insights:
(1) Enhanced Sensitivity. AuscultaBase achieved a sensitivity of 80%, substantially higher than the 38% attained by human experts, though at the expense of a lower specificity (64% versus 100%). The high sensitivity of the model indicates a reduced risk of missing diseased cases, thereby enabling earlier interventions—an essential factor in CHD detection. In practice, a synergistic diagnostic approach that combines the high sensitivity of AuscultaBase with the high specificity of experienced clinicians may offer the best of both worlds, optimizing early detection while minimizing false positives.
(2) Superior Diagnostic Accuracy in Young Patients. Overall, AuscultaBase achieved an accuracy of 72%, slightly surpassing the 69% accuracy of human experts. Notably, the model demonstrated marked superiority in younger age groups—achieving 77% accuracy in children under one year old compared to 58% for human experts, and 80% accuracy in children aged one to three years versus 60% for human experts. Early and accurate diagnosis in these critical age groups is vital; for example, timely intervention can prevent the progression of CHD-related complications such as pulmonary arterial hypertension (PAH) and subsequent Eisenmenger syndrome.
(3) Promising Role as a Diagnostic Assistant. The complementary strengths of AuscultaBase and human experts suggest a promising pathway toward integrated diagnostic systems. By leveraging the model’s high sensitivity to flag potential cases and utilizing the expertise of clinicians for confirmation, the overall diagnostic process can be both more accurate and more efficient. This integrative approach could transform clinical workflows in pediatric cardiology, ultimately leading to better patient outcomes.
In summary, our experimental results demonstrate that AuscultaBase not only outperforms existing baseline models across a variety of tasks but also shows significant promise as a clinical diagnostic tool. Its robust performance in the face of data imbalance, adaptability across diverse sound types and task formulations, and its potential to complement human expertise highlight its substantial potential to improve early detection and treatment in healthcare settings.
3 Discussion
This work presents a large-scale body sound corpus (AuscultaCorpus), an auscultation acoustic foundation model (AuscultaBase), and a comprehensive benchmark for evaluation (AuscultaBench). Our work addresses existing limitations in traditional auscultation practices through the integration of AI. By training on a diverse collection of body sounds using contrastive learning, our model shows a substantial improvement over existing pre-trained acoustic models, especially in imbalanced and limited-data settings. The full fine-tuning results reveal the robustness of our model in both binary and multi-class classification tasks, underscoring its potential for real-world clinical applications. The comparison of the auscultation efficacy between human experts and our model also indicates our model’s significance and clinical potential as a diagnostic assistant. This work emphasizes the value of combining traditional auscultation with AI-driven methods to deliver more reliable and accessible healthcare solutions, paving the way for the broader adoption of AI in medical diagnostics.
Given the cirtical need for rapid and cost-effective screening solutions, in this study, we proposed a comprehensive framework, including the AuscultaCorpus (a large-scale body sound corpus), the AuscultaBase (an auscultation acoustic foundation model), and the AuscultaBench (a comprehensive evaluation benchmark), to advance AI-assisted auscultation and improve the diagnostic capabilities of primary care providers while enabling scalable, high-quality body sound analysis. These contributions not only address the limitations of traditional auscultation practice, but also provide strong support for the application of AI in medical diagnosis.
The AuscultaCorpus offers a diverse set of body sound data, enabling the AuscultaBase model to learn more robust feature representations. This, in turn, improves performance in diverse diagnostic downstream tasks, outperforming baseline models. Notably, the model’s success validates AuscultaBase as a valuable tool for clinical practice. The AuscultaBench provides a structured framework for evaluating model performance across 16 tasks within two key diagnostic categories: abnormality detection and disease diagnosis. This ensures consistent, reproducible comparisons and drives progress in automated diagnostic methods.
AuscultaBase demonstrated strong generalization, achieving the highest Macro-F1 performance in 12 out of 16 tasks and leading in Micro-F1 for 10 tasks. This robust performance highlights the model’s versatility across different auscultation tasks. By contrast, baseline models such as OPERA-CT, AudioMAE, CLAP, and PANN lagged behind in terms of task completion and accuracy. Additionally, AuscultaBase showed excellent balanced accuracy, particularly in handling imbalanced datasets (refering to the Supplementary Information file), which is crucial for real-world clinical situations. In binary classification, it surpassed baseline models in terms of ROC curves and AUC value, further emphasizing its potential for clinical screening and early disease intervention.
Multi-dimensional evaluations also demonstrated AuscultaBase’s adaptability across various acoustic environments, confirming its ability to perform well in diverse real-world clinical settings. AuscultaBase not only has the ability to perform nuanced clinical assessments, but also has the ability to impact patient care through more precise disease identification. This robust performance in different acoustic environments demonstrates the adaptability of AuscultaBase, making it a versatile tool for encountering multiple sound sources and conditions in real-world clinical applications. This versatility suggests that AuscultaBase is able to effectively generalize to different problem formulations, which is an important quality for its application in different diagnostic scenarios.
In addition, in a comparative analysis with human experts, we further validated the clinical application prospects of AuscultaBase. In the comparison, AuscultaBase demonstrated high sensitivity, indicating a reduced risk of missed cases, thus enabling early intervention. Notably, the model showed a clear advantage in disease diagnosis accuracy in young patients, where early and accurate diagnosis is crucial for such critical age groups. AuscultaBase not only outperformed existing baseline models on a variety of tasks, but also showed significant application potential as a clinical diagnostic tool. Both the adaptability across different sound types and task formulations, and the potential to complement human expertise, highlight its great potential for improving early detection and treatment in healthcare settings.
However, some limitations may still remain in our study. Although AuscultaCorpus is large, it may not encompass all possible body sound types, which could lead to suboptimal performance when encountering rare or previously unknown sounds. Furthermore, while AuscultaBase has shown excellent performance across multiple tasks, its application in real-world clinical settings requires further validation. Additionally, the interpretability of the model needs improvement, so clinicians can better understand and trust its diagnostic results. This omission stems from the practical challenges associated with clinicians annotating precise murmur timestamps for each cardiac cycle, which consequently precluded our ability to conduct interpretability analysis on temporal prediction accuracy.
In view of the existing limitations, future research can be deepened from the following dimensions: First, continue to expand and diversify the corpus of body sound data to improve the generalization ability of the model. This includes the introduction of new acquisition tools such as smartphones and wearable devices to capture more comprehensive sound frequencies and background noises and enrich data diversity. Second, strengthen the validation and optimization of AuscultaBase in clinical settings. Through multi-center clinical trials, the performance and applicability of the model are evaluated in different medical institutions and regions. Close collaboration with clinicians is essential to jointly design trial plans, collect real clinical data, and provide timely feedback on model performance to iteratively optimize model parameters and improve diagnostic accuracy and reliability.
In conclusion, the proposed framework has great potential in the field of AI-assisted auscultation. The strong generalization ability, robustness, balanced accuracy and discrimination ability of AuscultaBase provide solid support for its wide application in clinical practice. Although our framework still has some limitations, in the future, we can continue to focus on corpus expansion and clinical validation to ensure the wide application and sustainable development of AI in medical diagnosis.
4 Online Methods
4.1 AuscultaCorpus: Pretraining Datasets
A total of 11 datasets containing cardiac, respiratory, and bowel sounds were collected to construct AuscultaCorpus, which consists of more than 40,000 recordings and 300 hours. Of the body sound datasets listed in Table 2, most datasets are publicly available, except for HSDreport and XHheartSound. The latter is a private dataset collected by Anonymous Hospital and has not yet been released as open source. The recordings were captured using various digital stethoscope versions positioned on the chest, heart, and abdomen. This diverse collection enhances the model’s ability to analyze body sound frequency and rhythm, ultimately improving its clinical auscultation applications. Additionally, a subset of labeled samples is held out from the pretraining dataset for subsequent downstream benchmarking. Please refer to the Supplementary Information file for more details about the data construction, auscultation positions, and patients’ metadata.